I wasn't sure where to post this question, so here goes.
The daily low temperatures in degrees Fahrenheit for a certain week were –3, 1, 4, –5, –10. What is the range of low temperatures for that week?
I understand range to be finding the difference between the highest and lowest numbers. The answer would be the range. But the multiple choice answers are 12, 13, 14, or 14 degrees.



Found 2 solutions by jim_thompson5910, checkley75:
Answer by jim_thompson5910(35256)

Range= highest number-lowest number
Range=(4)-(-10)

Range=(4)+(10)=14

So the range of the low temperatures is 14 degrees.

Answer by checkley75(3666)

THE DIFFERENCE BETWEEN THE HIGHEST (4) & THE LOWEST (-10)IS:
4-(-10)=14 FOR THE RANGE OF TEMPERATURES.
AN EASY WAY TO SOLVE THESE TYPE OF PROBLEMS IS TO DRAW A LINE. MARK THE CENTER OF THE LINE 0 AND MARK OF +10 UNITES & -10 UNITS. NOW PLOT THE POINTS +4 & -10. NOW COUNT ALL THE UNITS BETWEEN THESE 2 POINTS. YOU SHOULD GET 14 UNITS.